# The Hot Jupiter Radius Anomaly and Stellar Connections

### v1

@@ 3.2 Evolution Models
For planets consisting entirely of adiabatic and/or isothermal layers with a continuous temperature profile, the thermal state of the planet is reduced to a single variable. Typically this is the specific entropy of the envelope . A planet cools by releasing heat from the interior through its atmosphere, which depends on the thickness of the atmosphere (and therefore the mean molecular weight and gravity), its composition, and the incident flux from the star – highly irradiated atmospheres release heat from the interior more slowly. This can be specified as a total luminosity or an “intrinsic temperature” , which are related via , where is the Stefan-Boltzmann constant and R is the planet radius, usually at the radiative-convective boundary. Calculating the rate at which heat is released requires an atmosphere model that allows for both convective and radiative heat transport within the atmosphere down to the (RCB). This procedure is outside the scope of this chapter, but readers are instead directed to (Fortney et al. 2007) and the chapter of this book concerning radiative transfer. The rate of change of total energy in the planet can be written as
where is the hot Jupiter anomalous heating power (see Sec. 6.1 for estimates), is the heat generated by in the planet, is the release of tidal energy within the planet. This tidal heating could occur due to the damping of orbital eccentricity or obliquity between the planets rotational and orbital axes (Mardling 2007; Jackson et al. 2008, e.g.,); however, while it is likely important in certain cases (Millholland et al. 2020) it is not thought to explain the hot Jupiter radius anomaly on its own (Leconte et al. 2010), due to being short-lived on hot Jupiter orbits without an additional planet exciting eccentricity in the long term.
@@
Here, is the specific heat of the core, is the core mass, is the core temperature (as in this case we have assumed it is isothermal), is the total mass of the planet, indicates mass contained within a shell at the given point in the planet, and is the temperature as a function of pressure and specific entropy . With Eqs. 7 and 8, one can calculate the rate of change of the specific entropy as the product . This may be integrated forward using a standard ODE integrator to obtain the of the planet. It’s worth noting that for hot Jupiters these equations can be somewhat “stiff” (an ODE term outside the scope of this review), so it can be helpful to select an integrator that tolerates stiff ODEs such as backward differentiation formula (BDF) integrators.
@@ 4 Heating Depth and Reinflation
An important factor in how gaseous planets respond to heating is the depth at which the heat is deposited. This is because convection can only move heat upwards – hot fluid is less dense and rises through a colder medium, assuming a uniform composition. As such, heat injected near the planet’s surface can only slow cooling, not reverse it (e.g., Wu and Lithwick 2013, for the Ohmic dissipation case). Komacek and Youdin (2017) goes further, showing that heat must be injected at least at the (RCB) in order to even slow the cooling; heat deposited above this point will be re-radiated out too quickly to affect the planet’s radius or thermal evolution significantly. While originally the RCB was thought to be very deep in the planet (Guillot et al. 1996), accounting for the planets’ hot interiors indicates that the RCB will be at around 10 bar for most hot Jupiters (Thorngren et al. 2019a), moving to lower pressures with increasing metallicity and/or decreasing surface gravity.
A prediction of any model based on heating near the surface of the planet (see Sec. 5) is therefore that planets will cool very slowly, but cannot grow larger ( ) if the flux is later increased (e.g., Wu and Lithwick 2013; Komacek et al. 2020). Lopez and Fortney (2016) observes that this is testable by examining whether giant planets around evolved stars are disproportionately larger as a result of the increase in stellar luminosity. Grunblatt et al. (2016, 2017) finds two such giant planets and observes that they are slightly larger than typical hot Jupiters would be without reinflation, though it would be helpful to find additional cases to verify the trend and potentially measure the rate at which reinflation occurs.
@@ 5 Proposed Inflation Mechanisms
One of the earliest proposed explanations for the inflated radii of hot Jupiters was that the model atmosphere was significantly lower than in real hot Jupiters (Burrows et al. 2007, 2008). Such an error would cause models to cool too quickly, resulting in model planets that were too small. The advantages of this explanation include its lack of need for any new physics or a source of deep heating. However, it has a difficult time explaining the extremely large radii (up to ) that were subsequently discovered. As such, although this may be an important source of theoretical uncertainty in hot Jupiter atmosphere models, it does not appear to be the primary driver of their anomalous radii.
@@
Another solution suggested early on was that of (Bodenheimer et al. 2001, 2003). Here, tides raised on a planet by the parent star would damp out the planet’s eccentricity and spin-orbit obliquity, depositing the energy reduction into the planet’s interior as heat. However, this is a finite reservoir of energy and the tidal timescales for hot Jupiters are in general quite short (Jackson et al. 2008), so these processes would only be important in the earliest stages of a hot Jupiter’s life. Additionally, tidal forces scale strongly with semimajor axis and radius, not the incident flux as has been observed. As such, while this may be important for some planets (Millholland and Spalding 2020) (not just hot Jupiters), it is not favored as a general solution (Leconte et al. 2010; Thorngren and Fortney 2018). On the other hand, Arras and Socrates (2009) suggest that tidal heating may not arise from eccentricity or obliquity, but rather the tidal pull on the planet’s thermal bulge (Socrates 2013) which is offset from the sub-stellar point by winds. In this model, energy is gradually pulled from the (enormous) reservoir of the planet’s orbit and fed into the deep interior. A key benefit of this model is that the deep heating could potentially reinflate the planet. However, follow-up studies differ as to whether the scaling is able to match the correlation between radius and flux (Sarkis et al. 2020; Thorngren and Fortney 2018).
@@
Several proposed explanations for radius inflation can be broadly categorized together as fluid dynamic in nature. These share the common theme of arguing that the existing 1D atmospheric models are missing important physics that change the temperature-pressure profile of the deep atmosphere and suppress the release of heat from the interior. One such model is the Youdin and Mitchell (2010) proposal, in which the high-speed winds of the upper atmosphere generate turbulence beneath, dragging hot gas downward against the buoyancy force, producing a hotter and deeper RCB. Such an effect would slow the cooling process of the hot Jupiter substantially, though it would not by itself be capable of reinflating planets that had already cooled.
@@
More recently, Tremblin et al. (2017) proposed that the inflation effect may be resolved in higher-dimensional atmosphere models. These 3D global circulation models (GCMs) are computationally expensive, especially when modeling the deep atmosphere where thermal timescales can be very long. Instead, they solve for the steady-state solutions in 2D at various latitudes. They find that longitudinally around the planet (steady-state flows, rather than turbulence) has the effect of burying heat deeper in the atmosphere than is indicated by 1D models. The result is suppressed cooling, hotter interiors, and therefore substantially larger radii. In fact, their model predictions were somewhat larger than the observed radii – they suggest among other things that this may be the result of the 2D model that will not occur in a full 3D model. Sainsbury-Martinez et al. (2019) studied this effect in 3D and supported their findings, suggesting that future studies use substantially longer integration times and initialize their simulations with temperature-pressure profiles above equilibrium for faster convergence.
@@
A somewhat related proposal to the fluid dynamical models relies on a common feature of GCM results for hot Jupiters: a fast eastward (super-rotating) jet stream (e.g., Showman and Polvani 2011) carrying gas around the tidally locked planet. In the presence of a planetary magnetic field, this jet may cause ions (especially Na) and free electrons to drift such that a current flows through the atmosphere from the polar regions to the equator (Batygin and Stevenson 2010). This current flows back to the equator through the deeper parts of the atmosphere, depositing energy as a result of the material’s electrical resistivity. This process, called , has been proposed as a source of the heating required to inflate hot Jupiters (Batygin and Stevenson 2010; Batygin et al. 2011).
@@
Originally it was thought that this would result in ultra-hot Jupiters inflating to the point that they overflow their Roche lobes, but subsequent studies disagree (Spiegel and Burrows 2013); Menou (2012); Rauscher and Menou (2013) found that magnetic braking can significantly reduce the velocity of the atmospheric winds, in turn reducing the effectiveness of Ohmic dissipation. The peak efficiency of turning incident light energy into interior heating was found to occur at K by Menou (2012) and K by Rogers and Komacek (2014). Additionally, it is worth noting that this heating mechanism is fairly shallow, perhaps around 10 bars (Rogers and Komacek 2014), so it is not capable of causing reinflation (Wu and Lithwick 2013, see Sec. 4). Pu and Valencia (2017) investigated the process for hot Neptunes and found that Ohmic dissipation may operate in a similar manner in those objects as in hot Jupiters. On the other hand, Huang and Cumming (2012) modeled the heating from Ohmic dissipation and instead argue that the total heating is not sufficient to explain the radii observed.
@@ 6 Observable Effects
In this section I’ll review some of the observation-based constraints that have been made to the nature of the hot Jupiter radius anomaly. Because of the extremely large number of diverse observations made of hot Jupiters, I’ll leave to other chapters observations which do not directly relate to the inflationary mechanism, especially transmission and emission spectra (discussed elsewhere in this book).
@@ 6.1 Population Trends in Radius
Given the large and growing population of observed hot Jupiters, a helpful approach to understanding them has been to consider their properties as a population, averaging out individual variations in metallicity or observational errors. Laughlin et al. (2011) did this by calculating the radius anomaly , the radius minus the expected radius for an evolved metal-free planet, for the 90 transiting hot Jupiters known at the time. They found that the anomaly was strongly related to the effective temperature, with , and argue that this appears consistent with the Ohmic dissipation model. While he supports the observation of a correlation between radius anomaly and incident flux, a key result is that the heating is not a fixed fraction of the incident flux, but one that itself increases with flux.
Following a similar thread, Thorngren and Fortney (2018) analyzed the then 318 hot Jupiters with masses and radii, directly including the internal heating power as a structure model parameter. The degeneracy between heating and composition was resolved by imposing the Thorngren et al. (2016) mass-metallicity relation as a prior. For individual planets, this doesn’t resolve the degeneracy fully, but applied to a large population it becomes possible to measure heating trends. They found that the as a fraction of flux rose to a peak of about 2.5% at 1570 K ( erg s-1 cm-1) and then declines to below 1% for the ultra-hot ( K) planets. They found a good fit using a Gaussian in log-flux to predict the heating:
@@
Sarkis et al. (2020) conducted a similar study, seeking to extract the heating efficiency from the population of hot Jupiters with a Bayesian hierarchical model. Unlike Thorngren and Fortney (2018), they included a number of rigorous checks including whether they could retrieve the modeled value from a synthesized population (they did), checks for prior sensitivity (their results were insensitive), and more extensive comparisons with predicted heating from various models. They too found a peak in heating efficiency of about 2.5%, but at a higher temperature of K; this difference does not appear to be due to choice of prior or the statistical framework, but may be a difference in atmosphere models. They present several parameterizations for estimating , but we’ll just reproduce the Gaussian here (in terms of ) for reference and comparison:
@@
where is the Stefan-Boltzmann constant. Remember that depends on flux (or ), so this is not to be understood as a linear relationship between and . Still, using one of the formulas for allows one to calculate the relatively easily (e.g., Thorngren and Fortney 2019), which can be particularly helpful for studies where an estimate of it is needed; we’ll discuss such applications in the next section.
@@ 6.2 Intrinsic Temperature and Radiative-Convective Boundary
A hot interior has the potential to substantially affect the structure and heat flow through the planet’s atmosphere. The (see also Sec. 3.2) is a way to describe the energy entering the atmosphere from below. It participates in the heat balance of the atmosphere as:
@@
The intrinsic temperature can have detectable impacts on the observable atmosphere (see Fortney et al. 2020; Gao et al. 2020); for example Sing et al. (2019a) argue that the presence of Fe and Mg in the atmosphere of WASP-121 (which they observed with HST) requires a of at least 500 K to avoid the atoms condensing in the atmosphere too deep to be observed. This is consistent with the deep heating case – Sarkis et al. (2020) calculate a K for that planet. The presence of absence of these cold traps is therefore a useful diagnostic of the intrinsic temperature of some planets (Parmentier et al. 2016; Fortney et al. 2020). The atmospheric chemistry may also be affected by , as higher values result in a higher , pushing the chemical abundances further from equilibrium. The intrinsic temperature is also an important input in global circulation models (GCMs) of giant planets (May et al. 2021; Steinrueck et al. 2021; Lee et al. 2022, e.g.,), though in that field it is referred to as internal temperature.
The intrinsic temperature and equilibrium temperature also directly affect the location of the (RCB, see Sec. 4): a higher pushes the RCB to higher pressures, whereas a high pushes it to lower pressures (e.g., Guillot et al. 1996; Fortney et al. 2007; Thorngren and Fortney 2019). When hot Jupiters were initially discovered, it was believed that their very high equilibrium temperatures would move the RCB very deep into the planet, at around 1 kilobar (e.g., Guillot et al. 1996). However, these models were based on RV discoveries, which had not yet revealed the radius anomaly problem. If large radii indeed imply large intrinsic temperatures, Thorngren et al. (2019a) and Sarkis et al. (2020) both find that the RCB is pushed back up to 1-100 bars, depending on the planets’ exact equilibrium temperature and surface gravity. As Sarkis et al. (2020) note, heat must be deposited below the RCB to have a significant impact on the radius (Komacek et al. 2017), so a shallower RCB imposes less stringent requirements on the pressures of heat deposition. They argue further that the same argument could be applied to the temperature advection mechanism of Tremblin et al. (2017).

### v3

@@ 3.2 Evolution Models
For planets consisting entirely of adiabatic and/or isothermal layers with a continuous temperature profile, the thermal state of the planet is reduced to a single variable. Typically this is the specific entropy of the envelope . A planet cools by releasing heat from the interior through its atmosphere, which depends on the thickness of the atmosphere (and therefore the mean molecular weight and gravity), its composition, and the incident flux from the star – highly irradiated atmospheres release heat from the interior more slowly. This can be specified as a total luminosity or an “intrinsic temperature” , which are related via , where is the Stefan-Boltzmann constant and R is the planet radius, usually at the radiative-convective boundary. Calculating the rate at which heat is released requires an atmosphere model that allows for both convective and radiative heat transport within the atmosphere down to the radiative-convective boundary (RCB). This procedure is outside the scope of this chapter, but readers are instead directed to (Fortney et al. 2007) and the chapter of this book concerning radiative transfer. The rate of change of total energy in the planet can be written as
where is the hot Jupiter anomalous heating power (see Sec. 6.1 for estimates), is the heat generated by radioactive decays in the planet, is the release of tidal energy within the planet. This tidal heating could occur due to the damping of orbital eccentricity or obliquity between the planets rotational and orbital axes (Mardling 2007; Jackson et al. 2008, e.g.,), decaying to zero as the orbit circularizes and the rotation aligns to the orbit.
@@
Here, is the specific heat of the core, is the core mass, is the core temperature (as in this case it is assumed to be isothermal), is the total mass of the planet, indicates mass contained within a shell at the given point in the planet, and is the temperature as a function of pressure and specific entropy . With Eqs. 7 and 8, one can calculate the rate of change of the specific entropy as the product . This may be integrated forward using a standard ODE (ordinary differential equation) integrator to obtain the thermal evolution of the planet. It’s worth noting that for hot Jupiters these equations can be somewhat “stiff” (an ODE term outside the scope of this review), so it can be helpful to select an integrator that tolerates stiff ODEs such as backward differentiation formula (BDF) integrators. Fig. 4 shows the radius evolution of a Jupiter-mass planet following this method for various assumptions about the anomalous heating and intrinsic temperature.
@@ 4 Heating Depth and Reinflation
An important factor in how gaseous planets respond to heating is the depth at which the heat is deposited. This is because convection can only move heat upwards – hot fluid is less dense and rises through a colder medium, assuming a uniform composition. As such, heat injected near the planet’s surface can only slow cooling, not reverse it (e.g., Wu and Lithwick 2013, for the Ohmic dissipation case). Komacek and Youdin (2017) go further, showing that heat must be injected at least at the radiative-convective boundary (RCB) in order to even slow the cooling; heat deposited above this point will be re-radiated out too quickly to affect the planet’s radius or thermal evolution significantly. While originally the RCB was thought to be very deep in the planet (Guillot et al. 1996), accounting for the planets’ hot interiors indicates that the RCB will be at around 10 bar for most hot Jupiters (Thorngren et al. 2019a), moving to lower pressures with increasing metallicity and/or decreasing surface gravity.
A prediction of any model based on heating near the surface of the planet (see Sec. 5) is therefore that planets will cool very slowly, but cannot grow larger (reinflate) if the flux is later increased (e.g., Wu and Lithwick 2013; Komacek et al. 2020). Lopez and Fortney (2016) observe that this is testable by examining whether giant planets around evolved stars are disproportionately larger as a result of the increase in stellar luminosity. Grunblatt et al. (2016, 2017) find two such giant planets and observes that they are slightly larger than typical hot Jupiters would be without reinflation, though it would be helpful to find additional cases to verify the trend and potentially measure the rate at which reinflation occurs.
@@ 5 Proposed Inflation Mechanisms
One of the earliest proposed explanations for the inflated radii of hot Jupiters was that the model atmosphere opacity was significantly lower than in real hot Jupiters (Burrows et al. 2007, 2008). Such an error would cause models to cool too quickly, resulting in model planets that were too small. The advantages of this explanation include its lack of need for any new physics or a source of deep heating. However, it has a difficult time explaining the extremely large radii (up to ) that were subsequently discovered. As such, although this may be an important source of theoretical uncertainty in hot Jupiter atmosphere models, it does not appear to be the primary driver of their anomalous radii.
@@
Another solution suggested early on was that of tidal heating (Bodenheimer et al. 2001, 2003). Here, tides raised on a planet by the parent star would damp out the planet’s eccentricity and spin-orbit obliquity, depositing the energy reduction into the planet’s interior as heat. However, this is a finite reservoir of energy and the tidal timescales for hot Jupiters are in general quite short (Jackson et al. 2008), so these processes would only be important in the earliest stages of a hot Jupiter’s life. Additionally, tidal forces scale strongly with semimajor axis and radius, not the incident flux as has been observed. As such, while this may be important for some planets (Millholland et al. 2020) (not just hot Jupiters), it is not favored as a general solution (Leconte et al. 2010; Thorngren and Fortney 2018). On the other hand, Arras and Socrates (2009) suggest that tidal heating may not arise from eccentricity or obliquity, but rather the tidal pull on the planet’s thermal bulge (Socrates 2013) which is offset from the sub-stellar point by winds. In this thermal tides model, energy is gradually pulled from the (enormous) reservoir of the planet’s orbit and fed into the deep interior. A key benefit of this model is that the deep heating could potentially reinflate the planet. However, follow-up studies differ as to whether the scaling is able to match the correlation between radius and flux (Sarkis et al. 2020; Thorngren and Fortney 2018).
@@
Several proposed explanations for radius inflation can be broadly categorized together as fluid dynamic in nature. These share the common theme of arguing that the existing 1D atmospheric models are missing important physics that change the temperature-pressure profile of the deep atmosphere and suppress the release of heat from the interior. One such model is the mechanical greenhouse Youdin and Mitchell (2010) proposal, in which the high-speed winds of the upper atmosphere generate turbulence beneath, dragging hot gas downward against the buoyancy force, producing a hotter and deeper RCB. Such an effect would slow the cooling process of the hot Jupiter substantially, though it would not by itself be capable of reinflating planets that had already cooled.
@@
More recently, Tremblin et al. (2017) proposed that the inflation effect may be resolved in higher-dimensional atmosphere models. These 3D global circulation models (GCMs) are computationally expensive, especially when modeling the deep atmosphere where thermal timescales can be very long. Instead, they solve for the steady-state solutions in 2D at various latitudes. They find that temperature advection longitudinally around the planet (steady-state flows, rather than turbulence) has the effect of burying heat deeper in the atmosphere than is indicated by 1D models. The result is suppressed cooling, hotter interiors, and therefore substantially larger radii. In fact, their model predictions were somewhat larger than the observed radii – they suggest among other things that this may be the result of the 2D model that will not occur in a full 3D model. Sainsbury-Martinez et al. (2019) studied this effect in 3D and supported their findings, suggesting that future studies use substantially longer integration times and initialize their simulations with temperature-pressure profiles above equilibrium for faster convergence.
@@
A somewhat related proposal to the fluid dynamical models relies on a common feature of GCM results for hot Jupiters: a fast eastward (super-rotating) jet stream (e.g., Showman and Polvani 2011) carrying gas around the tidally locked planet. In the presence of a planetary magnetic field, this jet may cause ions (especially Na) and free electrons to drift such that a current flows through the atmosphere from the polar regions to the equator (Batygin and Stevenson 2010). This current flows back to the poles through the deeper parts of the atmosphere, depositing energy as a result of the material’s electrical resistivity. This process, called Ohmic dissipation, has been proposed as a source of the heating required to inflate hot Jupiters (Batygin and Stevenson 2010; Batygin et al. 2011).
@@
Originally it was thought that this would result in ultra-hot Jupiters inflating to the point that they overflow their Roche lobes, but subsequent studies disagree (Spiegel and Burrows 2013); Menou (2012); Rauscher and Menou (2013) found that magnetic braking can significantly reduce the velocity of the atmospheric winds, in turn reducing the effectiveness of Ohmic dissipation. The peak efficiency of turning incident light energy into interior heating was found to occur at K by Menou (2012) and K by Rogers and Komacek (2014). Additionally, it is worth noting that this heating mechanism is fairly shallow, perhaps around 10 bars (Rogers and Komacek 2014), so it is not capable of causing reinflation (Wu and Lithwick 2013, see Sec. 4). Pu and Valencia (2017) investigated the process for hot Neptunes and found that Ohmic dissipation may operate in a similar manner in those objects as in hot Jupiters. On the other hand, Huang and Cumming (2012) modeled the heating from Ohmic dissipation and instead argue that the total heating is not sufficient to explain the radii observed among hot Jupiters.
@@ 6 Observable Effects
This section reviews some of the observation-based constraints that have been made to the nature of the hot Jupiter radius anomaly. Because of the extremely large number of diverse observations made of hot Jupiters, it leaves to other chapters observations which do not directly relate to the inflationary mechanism, especially transmission and emission spectra (discussed elsewhere in this book).
@@ 6.1 Population Trends in Radius
Given the large and growing population of observed hot Jupiters, a helpful approach to understanding them has been to consider their properties as a population, averaging out individual variations in metallicity or observational errors. Laughlin et al. (2011) did this by calculating the radius anomaly , the radius minus the expected radius for an evolved metal-free planet, for the 90 transiting hot Jupiters known at the time. They found that the anomaly was strongly related to the effective temperature, with , and argue that this appears consistent with the Ohmic dissipation model. While they support the observation of a correlation between radius anomaly and incident flux, a key result is that the heating is not a fixed fraction of the incident flux, but one that itself increases with flux.
Following a similar thread, Thorngren and Fortney (2018) analyzed the then 318 hot Jupiters with masses and radii, directly including the internal heating power as a structure model parameter. The degeneracy between heating and composition was resolved by imposing the Thorngren et al. (2016) mass-metallicity relation as a prior. For individual planets, this doesn’t resolve the degeneracy fully, but applied to a large population it becomes possible to measure heating trends. They found that the heating efficiency as a fraction of flux rose to a peak of about 2.5% at 1570 K ( erg s-1 cm-2) and then declines to below 1% for the ultra-hot ( K) planets. They found a good fit using a Gaussian in log-flux to predict the heating:
@@
Sarkis et al. (2020) conducted a similar study, seeking to extract the heating efficiency from the population of hot Jupiters with a Bayesian hierarchical model. Unlike Thorngren and Fortney (2018), they included a number of rigorous checks including whether they could retrieve the modeled value from a synthesized population (they did), checks for prior sensitivity (their results were insensitive), and more extensive comparisons with predicted heating from various models. They too found a peak in heating efficiency of about 2.5%, but at a higher temperature of K; this difference does not appear to be due to choice of prior or the statistical framework, but may be a difference in atmosphere models. They present several parameterizations for estimating ; here is the Gaussian model (in terms of ) for reference and comparison:
@@
where is the Stefan-Boltzmann constant. Remember that depends on flux (or ), so this is not to be understood as a linear relationship between and . Still, using one of the formulas for allows one to calculate the relatively easily (e.g., Thorngren and Fortney 2019), which can be particularly helpful for studies where an estimate of it is needed; such applications are discussed in the next section.
@@ 6.2 Intrinsic Temperature and Radiative-Convective Boundary
A hot interior has the potential to substantially affect the structure and heat flow through the planet’s atmosphere. The intrinsic temperature (see also Sec. 3.2) is a way to describe the energy entering the atmosphere from below. It participates in the heat balance of the atmosphere as:
@@
The intrinsic temperature can have detectable impacts on the observable atmosphere (see Fortney et al. 2020; Gao et al. 2020); for example Sing et al. (2019) argue that the presence of Fe and Mg in the atmosphere of WASP-121 (which they observed with HST) requires a of at least 500 K to avoid the atoms condensing in the atmosphere too deep to be observed. This is consistent with the deep heating case – Sarkis et al. (2020) calculate a K for that planet. The presence of absence of these cold traps is therefore a useful diagnostic of the intrinsic temperature of some planets (Parmentier et al. 2016; Fortney et al. 2020). The atmospheric chemistry may also be affected by , as higher values result in a higher eddy mixing parameter , pushing the chemical abundances further from equilibrium. The intrinsic temperature is also an important input in global circulation models (GCMs) of giant planets (May et al. 2021; Steinrueck et al. 2021; Lee et al. 2022, e.g.,), though in that field it is referred to as internal temperature.
The intrinsic temperature and equilibrium temperature also directly affect the location of the radiative-convective boundary (RCB, see Sec. 4): a higher pushes the RCB to higher pressures, whereas a high pushes it to lower pressures (e.g., Guillot et al. 1996; Fortney et al. 2007; Thorngren and Fortney 2019). When hot Jupiters were initially discovered, it was believed that their very high equilibrium temperatures would move the RCB very deep into the planet, at around 1 kilobar (e.g., Guillot et al. 1996). However, these models were based on RV discoveries, which had not yet revealed the radius anomaly problem. If large radii indeed imply large intrinsic temperatures, Thorngren et al. (2019a) and Sarkis et al. (2020) both find that the RCB is pushed back up to 1-100 bars, depending on the planets’ exact equilibrium temperature and surface gravity. As Sarkis et al. (2020) note, heat must be deposited below the RCB to have a significant impact on the radius (Komacek et al. 2017), so a shallower RCB imposes less stringent requirements on the pressures of heat deposition. They argue further that the same argument could be applied to the temperature advection mechanism of Tremblin et al. (2017).